QUANTUM TUNNELING IN THE MIDRANGE VIBRATIONAL FUNDAMENTALS OF TROPOLONE

\underline{RICHARD L. REDINGTON} and THERESA E. REDINGTON , Department of Chemistry and Biochemistry, Texas Tech University, Lubbock, TX 79409; ROBERT L. SAMS, Wiley Environmental Molecular Sciences Laboratory, Pacific Northwest National Laboratory, Richland, WA 99352.

The FTIR spectrum of tropolone(OH) vapor in the 1175 - 1700 cm^{−1} region is reported at 0.0025 and 0.10 cm^{−1} spectral resolutions. Estimates for the spectral doublet component separations for 11 fundamentals are based on the measurement of sharp Q branch peaks. The three highest of six a_{1} modes show splittings Δ_{v} ≈ Δ_{0} (the known ZP splitting value), and four of the five observed b_{2} modes show splittings Δ_{v} ≈ 0.90 Δ_{0}. The remaining four splittings fall 4 to 14 % from Δ_{0}, including Δ_{8} (nominal COH bending) = 1.10 Δ_{0}. The data show that significant anharmonic behavior accompanies the mild perturbations of midrange vibrational excitations on the ZP tunneling mechanism. The results suggest the onset of fast IVR processes that can delocalize and smooth the imprint of the vibrational excitation energy on the ZP PES topography before the tautomerization process can occur. Further, the results suggest dynamical characteristics of the b_{2} modes include a damping effect on the tunneling process.